OpenStudy (anonymous):

determine the values of x where the tangent line is horizontal for the function f(x)=1/((x^2-1)(x-4))

OpenStudy (anonymous):

\[f(x)= \frac{ 1 }{(x^2-1)(x-4) }\]

OpenStudy (anonymous):

@hilbertboy96

OpenStudy (anonymous):

ah these are always fun. Well tell me if you know when are tangent lines to a function horizontal? There is a special term we call them

OpenStudy (anonymous):

i forgot the name, but i know that its when the slope/derivative is =0 right?

OpenStudy (anonymous):

yes they are called critical points and that is when f'(x)=0

OpenStudy (anonymous):

so we need to find f'(x) and the simpilest way to do this will most likely be to actually expand that bottom then do a composition rule

OpenStudy (anonymous):

so the bottom would be \[x^3-4x^2-x+4\]

OpenStudy (anonymous):

correct and to make it easier we can rewrite the function as \[f(x) = (x^3-4x^2-x+4)^{-1}\] right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

now we can take the derivative which as i said is a composition function so we apply the extended power rule \[f'(x)=-(x^3-4x^2-x+4)^{-2}(3x^2-8x-1)\] rewrite it and we get \[\frac{ -3x^2+8x+1 }{(x^3-4x^2-x+4)^{2} }\] and to make f'(x) = 0 we set the top equal to zero even though critical points are when f'(x) = 0 or does not exist but the do not exist spots are just vertical asymptotes so \[0=-3x^2+8x-1\] Solve and you get your x-values

OpenStudy (anonymous):

find them yet?

OpenStudy (anonymous):

i got\[\frac{ -(\sqrt{13} -4)}{ 3 } and \frac{ \sqrt{13}+4 }{3 }\]

OpenStudy (anonymous):

and it says its wrong :/

OpenStudy (anonymous):

hmm rechecking work

OpenStudy (anonymous):

wait did you go off of my 0=-3x^2+8x-1?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

the -1 should be +1

OpenStudy (anonymous):

sorry lol

OpenStudy (anonymous):

oh okay, let me change it

OpenStudy (anonymous):

lol its okay

OpenStudy (anonymous):

got it :), thanks, it came out to be \[x=\frac{ -(\sqrt{19}-4) }{3 } and \frac{ \sqrt{19} +4}{ 3 }\]